# Critical states exhibit invariance in both position and momentum spaces

### v2

@@ Abstract
The critical states of disordered systems are intriguing subjects within the realm of condensed matter physics and complex systems. These states manifest in materials where disorder plays a significant role, and are distinguished by their multifractal structure and self-similarity. However, accurately characterizing critical states continues to pose a significant challenge. In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains. More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space. Subsequent numerical simulations validate the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
@@ I Introduction
The concept of critical states Critical in various physical systems pertains to the behaviors and characteristics exhibited during phase transitions or as they approach critical points. The examination of critical states provides profound insights into the nature of order, disorder, and collective behavior across diverse disciplines Sachdev1 . These states are fundamental for comprehending a wide range of phenomena in statistical physics Critical0 , condensed matter physics Critical1 , and complex systems Critical2 . As a defining feature of critical phenomena Critical3 , fluctuations in critical states become correlated over large distances, with the system demonstrating self-similar behavior at different scales. Critical fluctuations have been observed in various systems, including phase transitions in ferromagnets-antiferromagnets Critical4 , percolation thresholds in complex networks Critical5 , and photonic quasicrystals photonic . As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of complex systems.
Recently, significant advancements have been achieved in the investigation of critical states of quantum disordered systems, specifically focusing on multifractal critical wave functions Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 . Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging. Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and accurately characterizing them.
Bohr’s complementarity principle MQM provides us with inspiration to bridge this gap. According to the complementarity principle, it is impossible to simultaneously observe the wave nature and particle nature of matter under different representations, such as position space and momentum space. Matter exhibits different properties in these representations. The relationship between position space and momentum space is primarily established through the Fourier transform; position space serves as a mathematical framework for describing the location of particles or objects in physical space, while momentum space is perceived as the duality of position space QM . The critical state can be considered as a third state in addition to particle and wave properties, thus it can be reasonably inferred that the critical state may exhibit some invariance in both position space and momentum space.
@@
Based on this duality perception, Liu and Xia propose an explicit criterion Liu-Critical for precisely characterizing critical states. They assert that the Lyapunov exponents of critical states should simultaneously be 0 in both position space and momentum space. The criterion definitively distinguishes critical states from extended and localized states, as the Lyapunov exponents of the latter two are not equal in position space and momentum space. If one exponent is 0, then the dual exponent must be greater than 0. In physics, this also implies that extended and localized states cannot simultaneously exhibit delocalization in position space and momentum space, as illustrated in Fig. 1(a) and (b). In contrast, critical states should demonstrate delocalization, multifractality, and self-similarity in both position space and momentum space, as depicted in Fig. 1(c) and (d). These findings clearly illustrate that critical states maintain a certain kind of invariance across two dual spaces.
Starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number. Contrastively, so far we have not obtained a closed-form solution for the critical state, not even an approximate expression. Based on the delocalization, multifractality and self-similarity of critical states, we propose a hypothetical solution, that is, the critical state is a power-law modulated Weierstrass-like function,
@@
A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces? This study demonstrates that the invariance of the critical state is universal, encompassing commonly utilized observables in experiments such as the inverse participation ratio and information entropy.
@@ II Analytical derivation
The application of Lyapunov exponents to disordered systems Xia ; Jitomirskaya ; Avila ; Dwiputra focuses on calculating the decay rate of the wave function. For a given disordered potential, the conjectured solution for the wave function in the $n$-th lattice is $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$, where $\gamma\geq 0$ represents the Lyapunov exponent. A value of $\gamma=0$ indicates that $|\psi_{n}|\sim|\psi_{0}$, corresponding to an extended state; while a value of $\gamma>0$ indicates that $|\psi_{n}|\sim|\psi_{0}|e^{-\gamma n}$, corresponding to a localized state. This proposed solution provides a reasonable framework encompassing eigenstates of disordered systems and can be utilized to derive mathematical expressions for the Inverse Participation Ratio and information entropy with respect to $\gamma$.
The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 . A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
Another typical physical quantity, information entropy $S$, commonly known as Shannon entropy Shannon , is a metric for the unpredictability associated with a system. Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 ; IE1 . Utilized to quantify the level of unpredictability and uncertainty IE2 ; IE3 , a large $S$ indicates a high unpredictability, corresponding to extended states; while a low $S$ reflects a high predictability, corresponding to localized states.
@@
According to the Liu-Xia criterion Liu-Critical , critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. This can be mathematically expressed as the equality of the Lyapunov exponent of position space $\gamma$ and that of momentum space $\gamma_{m}$,
@@
Equation (5) represents the key innovation of this study, demonstrating the universal invariance of the critical state in both position space and momentum space. This concept applies not only to Lyapunov exponent, but also to other fundamental physical quantities such as IPR and $S$.
@@ III.1 Aubry-André-Harper model
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential at site $n$, and $E$ is the energy eigenvalue of the particle. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity. $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity). Thus, the potential $V_{n}$ at each lattice site $n$ is quasiperiodic because it is a non-repeating function with long-range correlations, unlike random disordered potentials that are purely random at each site.
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem model
where the meaning of parameters is the same as that of the AAH model. The QNE model [Eq. (9)] differs significantly from the AAH model [Eq. (6)], particularly on the right-hand side of the equation, making it no longer a linear eigenvalue problem like previous quasiperiodic systems. The QNE model holds a unique position in the field of Non-Hermitian physics NH1 ; NH2 ; NH3 ; NH4 ; NH5 ; NH6 ; NH7 , because non Hermitian can be introduced in three ways: through non-reciprocity in hopping terms, through complex momentum involving gain and loss of reciprocal hopping terms, or by introducing complex mass representing gain and loss of on-site potential. However, the QNE model does not fall into any of these three cases as it is not a linear eigenvalue problem.
@@
Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the AAH model, respectively. (d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the QNE model, respectively. Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. In the AAH model, IPR and $S$ are only equal between position space and momentum space at the phase transition point $V=2$, whereas in the QNE model, these physical quantities are equal within a large parameter range of $0<V\leq 2$.
@@ III.3 Verification results
We perform direct numerical diagonalization on the AAH model and QNE model, obtaining the eigenvalues and eigenstates separately. The total number of lattices in the system is set to 1000. For the AAH model, our focus is on the middle eigenstate; whereas for the QNE model, we concentrate on the eigenstate with an eigenvalue of 0.

### v3

# Exactly characterizing critical states: invariance in both position and momentum spaces

@@ Abstract
The critical state of disordered systems is an intriguing and popular topic in condensed matter physics, known for its non-ergodic properties such as multifractal structure and self-similarity. However, exactly characterizing critical states continues to pose a significant challenge so far. In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit invariance in both position and momentum spaces, leading to their delocalization in both dual spaces. Furthermore, keeping pace with the idea of Liu-Xia criterion, we demonstrate that physical quantities being directly observed in experiments, such as the inverse participation ratio and information entropy, also exhibit invariance in both position and momentum spaces. Subsequent numerical simulations verify the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
@@ I Introduction
Critical states, also known as multifractal states, in various physical systems Critical pertains to the characteristics exhibited by phase transitions or across critical points. The examination of critical states provides profound insights into the nature of order, disorder, and collective behavior Sachdev1 . These states are fundamental for comprehending a wide range of phenomena in statistical physics Critical0 , condensed matter physics Critical1 , and complex systems Critical2 . As a defining feature of critical phenomena Critical3 , fluctuations in critical states become correlated over large distances, with the system demonstrating self-similar behavior at different scales. Critical fluctuations have been observed in various systems, including phase transitions in ferromagnets-antiferromagnets Critical4 , percolation thresholds in complex networks Critical5 , and photonic quasicrystals photonic . As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of Anderson localization Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; zhou ; Biddle ; Zhang .
Recently, significant advancements have been achieved in the investigation of critical states in quantum disordered systems, specifically focusing on multifractal wave functions q1 ; q2 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg9 . Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging. Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and exactly characterizing them.
Bohr’s complementarity principle MQM provides us with inspiration to bridge this gap. According to the complementarity principle, it is impossible to simultaneously observe the wave nature and particle nature of matter under different representations, such as position (or real) space and momentum (or dual) space. The relation between position space and momentum space is primarily established through the Fourier transform; position space serves as a mathematical framework for describing the location of particles or objects in physical space, while momentum space is perceived as the duality of position space QM . Matter generally exhibits different localization properties in two mutually dual spaces. The critical state can be considered as a third state in addition to particle and wave properties, thus it can be reasonably inferred that the critical state may exhibit some invariance in both position space and momentum space.
@@
Based on this duality perception, Liu and Xia propose an explicit criterion Liu-Critical for precisely characterizing critical states, asserting that the Lyapunov exponents of critical states should simultaneously be 0 in both position and momentum space. The criterion definitively distinguishes critical states from extended and localized states, as the Lyapunov exponents of the latter two are not equal in position and momentum space. If one exponent is 0, then the dual exponent must be greater than 0. In more physical language, extended and localized states cannot simultaneously exhibit delocalization in position and momentum space, as illustrated in Fig. 1(a) and (b). In contrast, Liu-Xia criterion predicts that critical states should exhibit delocalization, multifractality, and self-similarity in both position and momentum space, as depicted in Fig. 1(c) and (d). Hence, starting from a strict mathematical formalization perspective, the Lyapunov exponent $\gamma$ in position space and $\gamma_{m}$ in momentum space can be utilized to exactly distinguish extended, localized and critical eigenstates in quantum disordered systems, which are summarized in Table 1. These findings clearly illustrate that critical states maintain a certain kind of invariance across two dual spaces.
Table 1: Classification of three typical eigenstates in quantum disordered systems according to Liu-Xia criterion.
Eigenstates Position Momentum Delocalization
Extended $\gamma=0$ $\gamma_{m}>0$ Position
Localized $\gamma>0$ $\gamma_{m}=0$ Momentum
Critical $\gamma=0$ $\gamma_{m}=0$ Both
Beyond those, starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number. Contrastively, so far we have not obtained a closed-form solution for the critical state, not even an approximate expression. Based on the delocalization, multifractality and self-similarity of critical states, we propose a hypothetical solution, that is, the critical state is a power-law modulated Weierstrass-like function,
@@
Unfortunately, the critical state has not been fully confirmed by experiments so far. A question arises: can we utilize Liu-Xia criterion to convincingly observe critical states in experiments? One shortcoming is that it is not possible to directly measure the Lyapunov exponent in experiments. Therefore, the properties of physical quantities that can be directly measured in experiments such as the inverse participation ratio and information entropy in position space and momentum space are the focus of this work. Our study demonstrates that the invariance of the critical state is universal, the physical quantities that can be directly observed in these experiments are also invariant in both position and momentum space.
@@ II Analytical derivation
The application of Lyapunov exponents to disordered systems Xia ; Jitomirskaya ; Avila ; Dwiputra focuses on calculating the decay rate of the wave function. For a given disordered potential, the conjectured solution for the wave function in the $n$-th lattice is $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$, where $\gamma\geq 0$ represents the Lyapunov exponent. A value of $\gamma=0$ indicates that $|\psi_{n}|\sim|\psi_{0}|$, corresponding to an extended state; while a value of $\gamma>0$ indicates that $|\psi_{n}|\sim|\psi_{0}|e^{-\gamma n}$, corresponding to a localized state. This proposed solution provides a reasonable framework encompassing eigenstates of disordered systems and can be utilized to derive mathematical expressions for the Inverse Participation Ratio and information entropy with respect to $\gamma$.
The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions gg1 ; gg2 . A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
Another typical physical quantity, information entropy $S$, commonly known as Shannon entropy Shannon , is a metric for the unpredictability associated with a system. Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 . Utilized to quantify the level of unpredictability and uncertainty IE2 ; IE3 , a large $S$ indicates a high unpredictability, corresponding to extended states; while a low $S$ reflects a high predictability, corresponding to localized states.
@@
According to the Liu-Xia criterion Liu-Critical , we have
@@
Equation (5) demonstrates the universal invariance of the critical state in both position space and momentum space. This concept applies not only to Lyapunov exponent, but also to other fundamental physical quantities such as IPR and $S$.
@@ III.1 Aubry-André-Harper model
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential at site $n$, and $E$ is the energy eigenvalue of the particle. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity. $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity). Thus, the potential $V_{n}$ is quasiperiodic because it is a non-repeating function with long-range correlations, unlike random disordered potentials that are purely random at each site.
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem model
where the meaning of parameters is the same as that of the AAH model. The QNE model [Eq. (9)] differs significantly from the AAH model [Eq. (6)], particularly on the right-hand side of the equation, making it no longer a linear eigenvalue problem like previous quasiperiodic systems. The QNE model holds a unique position in the field of Non-Hermitian physics NH1 ; NH3 ; NH4 ; NH5 ; NH6 ; NH7 , because non Hermitian can be introduced in three ways: through non-reciprocity in hopping terms, through complex momentum involving gain and loss of reciprocal hopping terms, or by introducing complex mass representing gain and loss of on-site potential. However, the QNE model does not fall into any of these three cases as it is not a linear eigenvalue problem.
@@
Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the middle eigenstate of AAH model, respectively. (d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the $E=0.4$ eigenstate of QNE model, respectively. Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. In the AAH model, IPR and $S$ are only equal between position space and momentum space at the phase transition point $V=2$, whereas in the QNE model, these physical quantities are equal within a large parameter range of $0<V\leq 2$.
@@ III.3 Verification results
We perform direct numerical diagonalization on the AAH model and QNE model, obtaining the eigenvalues and eigenstates separately. The total number of lattices in the system is set to 987. For the AAH model, our focus is on the middle eigenstate; whereas for the QNE model, we concentrate on the eigenstate with an eigenvalue of $E=0.4$.